OpenStudy (riley):

The recommended daily intake of thiamine (vitamin B1) is 1.5mg. Weighing 100g, a small apple contains 4x10^-2 mg of thiamine. How many of these small apples would you need to eat everyday to get your needed thiamine, if the apples were your sole source of the vitamin? O_O??

OpenStudy (anonymous):

10^-2 = 1/10^2 = 1/100 So you have 4 * 1/100 = 4/100 4/100 actually is 0.04 So you can divide 1.5 by 0.04 to figure out how many dumb apples you need 1.5 / 0.04 = 37.5 So eat 38 apples or take a vitamin

OpenStudy (anonymous):

each apple contains 4*10^-2 mg of thiamine so 1.5 / (4*10^-2) = 1.5 / 0.04 = 37.5 apples

OpenStudy (anonymous):

maybe you cant have half apple so 38

OpenStudy (riley):

thank you so much, both of you!!! showing the math like that really helped, thanks!!<3333